### v2

# Authoritarian Recursions: How Fiction, History, and AI Reinforce Control in Education, Warfare, and Discourse

Abstract
This article introduces the concept of authoritarian recursion to describe how artificial intelligence (AI) systems increasingly mediate control across education, warfare, and digital discourse. Drawing on critical discourse analysis and sociotechnical theory, the study reveals how AI-driven platforms delegate judgment to algorithmic processes, normalize opacity, and recursively reinforce behavioral norms under the guise of neutrality and optimization. Case studies include generative AI models in classroom surveillance, autonomous targeting in military AI systems, and content curation logics in platform governance.
Rather than treating these domains as disparate, the paper maps their structural convergence within recursive architectures of abstraction, surveillance, and classification. These feedback systems do not simply automate tasks—they encode modes of epistemic authority that disperse accountability while intensifying political asymmetries. Through cultural and policy analysis, the article argues that authoritarian recursion operates as a hybrid logic, fusing technical abstraction with state and market imperatives. The paper concludes by outlining implications for democratic legitimacy, human oversight, and the political design of AI governance frameworks.
This framework contributes to emerging debates on algorithmic accountability by foregrounding how recursion acts not merely as a technical function but as a sociopolitical instrument of control.
Keywords: authoritarian recursion, AI governance, algorithmic accountability, surveillance, recursive control, critical discourse analysis, sociotechnical systems, educational technology, military AI, platform regulation
Introduction
Artificial intelligence (AI) is often described in terms of promise—optimization, personalization, predictive precision—but it is equally a technology of recursion. It does not merely operate within existing systems; it loops back upon them, amplifying their logics, stabilizing their asymmetries, and obscuring their origins. This paper names this recursive mechanism authoritarian recursion: the systematic reinforcement and evolution of control structures through algorithmic automation.
Authoritarian recursion is not a metaphor. It describes how AI systems—deployed in warfare, education, and content governance—codify prior patterns of judgment, visibility, and exclusion. These systems are rarely neutral. Their architectures inherit militarized epistemologies, bureaucratic incentives, and market logics that prioritize efficiency over deliberation, surveillance over solidarity. What emerges is not a clean break with the past, but a patterned automation of its most coercive functions.
This argument unfolds across three interrelated sites: drone warfare, algorithmic proctoring, and content moderation. These domains are not chosen for breadth but for depth—they reveal how AI mediates moral judgment under conditions of scale, uncertainty, and institutional inertia. Far from discrete sectors, they form a recursive matrix in which power circulates and legitimates itself through code.
The analysis proceeds interpretively, drawing on critical media theory, relational ethics, and fictional narratives that illuminate the epistemic stakes of automation. Cultural texts like Black Mirror, The Terminator, and Nineteen Eighty-Four function not as speculative detours but as epistemological tools. They dramatize how automation, once delegated, escapes retraction—how decisions made by machines are increasingly insulated from human redress. As Cave and Dihal argue, such representations do not merely reflect public fears; they actively shape institutional responses to AI [1].
This inquiry also grounds itself in contemporary institutional discourse. Corporate white papers, governmental defense protocols, and educational AI policies offer insight into how algorithmic systems are rationalized and normalized. Reports such as Meta’s transparency roadmap [2], Google’s responsible AI framework [3], and the Global Network Initiative’s human rights standards [4] reveal not only what these actors claim AI can do, but also what they believe it ought to do—and for whom.
AI is not deployed in a political vacuum. It is embedded within governance architectures that frame ethical responsibilities, determine thresholds of acceptability, and allocate moral authority to technical systems. To understand this embedding is not to reject technology but to interrogate its premises: Who defines fairness? Who delegates oversight? Who becomes invisible under the regime of the algorithm?
This paper argues that authoritarian recursion is not an aberration but a structural tendency—a design pattern disguised as inevitability. Recognizing it demands we move beyond interface-level reforms and ask harder questions about institutional legitimacy, recursive design, and the politics of automation itself.
1 Literature Review
This section reviews the scholarly and historical terrain across three domains where AI technologies operationalize control: military automation, educational surveillance, and algorithmic curation in digital discourse. While each context has sector-specific concerns, they share structural features—namely, opacity, delegation of judgment, and recursive feedback—that embed normative authority into automated systems.
1.1 Military AI and the Automation of Violence
Artificial intelligence in military settings reveals a growing entanglement of automation and coercion. Systems such as semi-autonomous drones, predictive surveillance networks, and AI-assisted targeting platforms prioritize speed, precision, and operational efficiency—yet often at the expense of ethical deliberation and legal accountability. Marsili warns that “the removal of human decision-making from the use of lethal force creates a dangerous precedent,” undermining the very humanitarian principles that military law is intended to uphold [5].
These concerns align with the Fairness, Accountability, and Transparency (FAccT) framework, which argues that fairness in automated decision-making must extend beyond output metrics to include contextual sensitivity, procedural redress, and meaningful oversight [6, 7]. In military AI, accountability often becomes nominal—tethered to protocols rather than substantive ethical reflection.
Historical precedents further complicate the notion of technological neutrality. During World War II, IBM’s punch card infrastructure was deployed by Nazi Germany for logistics and census operations—systems later used to facilitate genocide [8]. These tools encoded forms of bureaucratic rationality eerily analogous to today’s algorithmic architectures. As Asaro argues, the distancing of lethal decisions from moral responsibility through automation represents a dangerous ethical shift [9].
Speculative fiction reinforces this critique. Narratives like The Terminator, RoboCop, and Black Mirror episodes offer dystopian imaginaries of militarized AI. According to Cave et al., “the future imagined in fiction is often realized not because it is inevitable, but because it is ideologically compatible with dominant institutions” [10, p. 75]. These cultural texts anticipate how automation narratives rationalize political authority and normalize autonomous violence.
1.2 Educational AI: Automation of Surveillance and Discipline
Educational technologies increasingly employ AI systems for purposes such as assessment, behavioral monitoring, and classroom management. One prominent application is automated proctoring software that uses facial detection, motion tracking, and audio analysis to identify potential cheating. While marketed as tools of academic integrity, these systems have been shown to "unfairly disadvantage students" with darker skin tones, especially Black students and women of color, due to algorithmic biases in face detection and flagging procedures [11]. Such tools routinely flag these students at significantly higher rates—up to six times more often—despite no evidence of increased cheating, raising critical concerns about surveillance, equity, and educational harm [11, 12].
Here, too, the FAccT triad is often invoked as a remedy. However, operational deployments rarely meet its normative thresholds. Mittelstadt et al. emphasize that fairness requires attention to context and historical inequalities, not just statistical parity [6]. Yet student-users typically lack access to the internal logic of these systems and have little recourse to challenge their outputs.
Noble argues that “algorithmic decision systems often act as new instruments of racial and economic profiling” [12, p. 34]. This echoes Selwyn’s concerns that digital surveillance in education creates a system “in which suspicion is automated and dissent is pathologized” [13]. Such platforms reproduce the logic of panoptic discipline, as theorized by Foucault, where constant surveillance internalizes conformity [14].
Moreover, the political values behind these technologies often go unquestioned. Gilliard and Selwyn contend that “continued adoption of proctoring technologies in public education exposes a fundamental clash of politics,” where commercial priorities of security and efficiency override pedagogical values of equity and trust [15, p. 197]. These systems operationalize discipline not through direct coercion, but through the automation of suspicion and reduction of students to behavioral data.
1.3 AI in Discourse and Propaganda: Curation as Control
Algorithmic curation now structures the informational environment of billions. AI systems deployed by platforms such as Facebook, YouTube, and TikTok optimize content delivery for engagement, not accuracy. Tufekci notes that such platforms “amplify divisive content by design, creating an infrastructure for affective polarization” [16]. This modulation of attention constitutes a new form of informational power.
While FAccT-based interventions such as algorithmic impact assessments or explainability mechanisms have been proposed for content governance, their efficacy remains limited by platform opacity and commercial disincentives. Mittelstadt et al. argue that transparency without enforceable accountability often reduces ethical AI to “ethical theatre” [6].
Gillespie emphasizes that platforms are not neutral hosts but “custodians of public discourse” who shape access to visibility through inscrutable recommendation logics [17, p. 197]. Zuboff describes this shift as “instrumentarian power,” wherein behavior is not repressed but tuned through predictive analytics and behavioral nudging [18, p. 377].
This mode of control parallels historical propaganda. The Nazi regime used print, film, and spectacle to synchronize public perception. Today’s algorithmic persuasion, however, operates at greater scale and granularity—executing individualized influence operations based on psychometric data and engagement profiles [19, 20]. Berardi calls this the “colonization of subjectivity,” wherein cognition itself becomes a site of commodification and control [21].
Synthesis: Toward a Unified Critique of AI Control
Across military, educational, and discursive sectors, artificial intelligence technologies consistently reinforce rather than disrupt authoritarian structures of governance. Despite their domain-specific implementations, these systems display convergent design logics and ethical risks. The comparative literature suggests that AI operates not merely as a tool but as a vector of normative reproduction—embedding and amplifying existing asymmetries of power, visibility, and voice.
Three interlocking patterns emerge consistently across the domains surveyed. First, AI systems introduce a profound opacity that severs decision-making from those it affects. Whether in autonomous weapons systems, algorithmic proctoring, or content recommendation engines, the logic of the algorithm is rendered inaccessible, both technically and institutionally. This opacity undermines the possibility of contestation and erodes the conditions necessary for democratic oversight. Second, intelligent systems displace relational judgment by translating moral decisions into statistical approximations. This delegation of judgment dehumanizes its subjects: individuals are abstracted into data points, and the contingent, situated nature of ethical discernment is flattened into binary outputs or risk profiles. Third, and most insidiously, these systems perpetuate normative drift. They inherit and amplify structural biases—racial, economic, epistemic—under the rhetorical cover of objectivity or innovation. As they automate decision-making, they also automate exclusion, encoding historical inequalities into seemingly neutral infrastructures. These patterns do not reflect the malfunction of AI, but its core affordances within existing power regimes.
Together, these patterns constitute what may be termed an authoritarian recursion—a self-reinforcing cycle in which AI technologies encode, naturalize, and propagate control logics across domains. As behavior is shaped by algorithmic outputs, those outputs become future inputs, creating recursive systems that reinforce the very assumptions they were built upon.
This synthesis lays the foundation for the case study analysis that follows, which further illustrates how intelligent systems materialize these recursive dynamics in real-world governance structures.
2 Methodology
The analysis proceeds from an interpretive, critical-theoretical perspective, aimed at understanding how artificial intelligence (AI) systems participate in the reproduction of normative control across different social domains. Rather than offering empirical generalizations or causal inferences, the approach foregrounds how meaning is made, contested, and stabilized through discourse, design narratives, and institutional practice.
At the core of this inquiry lies Critical Discourse Analysis (CDA), particularly as theorized by Norman Fairclough. He contends that “discourse is both constitutive and constituted; it contributes to the shaping of social structures while being shaped by them” [22]. In other words, how we talk about AI—whether as intelligent, efficient, autonomous, or fair—feeds back into the very architectures we build. CDA enables a focus on these feedback mechanisms: where language meets infrastructure, and where discourse becomes design.
The analysis draws on three main types of material: theoretical literature, cultural narratives, and institutional documents. First, foundational texts in AI ethics, surveillance capitalism, and platform governance provide conceptual grounding. These include works by Zuboff, Coeckelbergh, and Gillespie, whose insights into autonomy, control, and datafication shape the framing of recursive algorithmic systems [18, 23, 17].
Second, cultural narratives—especially speculative fiction such as Orwell’s Nineteen Eighty-Four [24], James Cameron’s The Terminator [25], and episodes from Charlie Brooker’s anthology series Black Mirror [26]—are employed not as empirical data, but as heuristic provocations. These stories surface emerging ethical dilemmas in ways that formal models often overlook. As Cave and Dihal argue, cultural imaginaries help structure how societies anticipate and evaluate technologies before they fully arrive [1].
Third, the analysis incorporates public-facing and institutional documents, including educational policies on algorithmic proctoring [13], defense white papers on autonomous weapons systems [9], and corporate reports outlining the principles and practices of algorithmic content moderation. For instance, transparency initiatives like Meta’s system cards [2], Google’s Responsible AI framework [3], and the Global Network Initiative’s Principles on Freedom of Expression and Privacy [4] reveal how AI ethics is being operationalized at scale. These materials allow the study to trace not only how AI systems are designed and deployed, but how their legitimacy is framed through policy discourse and institutional self-regulation.
What holds these sources together is not empirical coherence but thematic resonance. All engage with questions of visibility, judgment, autonomy, and legitimacy—questions that lie at the center of both AI deployment and ethical inquiry. While some might ask whether these sources can be analyzed together, the real question is: how could they not be, given how power now flows across sectors and interfaces?
Fictional systems like Skynet, for example, are not invoked because they “predict” the future but because they illuminate a logic already at work: the automation of escalation, the erasure of moral deliberation, and the insulation of decision-making from relational accountability. Skynet is less a fantasy than a frame—one that helps articulate why human oversight cannot be treated as an afterthought in system design.
This interpretive approach has its limitations. It does not claim neutrality, nor does it offer quantifiable findings. But it does aim to clarify the stakes of a rapidly changing technological order by drawing lines between the past, the possible, and the political. The hope is that such lines might help us see AI systems not just as tools to be used, but as structures to be questioned.
3 Case Studies
AI does not act in isolation—it is embedded in systems that shape how knowledge, risk, and authority are organized. The following three domains—education, warfare, and discourse—reveal how intelligent systems delegate judgment, automate suspicion, and reinforce pre-existing norms through recursive feedback loops. These cases are not offered as an exhaustive taxonomy but as analytical vignettes into the institutional logics that AI both reflects and reconfigures.
3.1 Education: Algorithmic Proctoring and Epistemic Discipline
Remote proctoring platforms such as Proctorio and ExamSoft are marketed as scalable solutions for maintaining academic integrity in digital environments. Yet these systems pre-encode suspicion: gaze aversion, posture shifts, ambient silence—each is parsed as a probabilistic indicator of misconduct. In doing so, they invert the burden of proof and presume bad faith as a system default.
As Gilliard and Selwyn argue, such platforms enact “a shift in educational values, from those grounded in equity and social justice to those centered on control, surveillance, and suspicion” [15]. The proctor is no longer a relational figure but a predictive apparatus—opaque, unaccountable, and disembedded from context.
This dynamic echoes Foucault’s account of disciplinary power, where visibility enforces internalized control: “He who is subjected to a field of visibility… assumes responsibility for the constraints of power” [14, p. 202]. Algorithmic proctoring does not merely monitor students—it restructures the epistemology of assessment. Pedagogy becomes risk classification. Trust becomes anomaly detection. Judgment is replaced by metrics optimized for auditability. In this configuration, learning is subordinated to legibility.
3.2 Warfare: Autonomous Weapons and Delegated Lethality
Autonomous weapons systems (AWS), including platforms like Israel’s Harpy drone and the U.S. MQ-9 Reaper, are designed to identify and engage targets with minimal or no human oversight. Framed as innovations in precision warfare, they operationalize a deeper abstraction: the displacement of moral agency by machinic logics.
Peter Asaro warns that “removing humans from the process of selecting and engaging targets dehumanizes lethal decision-making and thereby undermines fundamental principles of human rights and international humanitarian law” [9, p. 696]. Crucially, this is not just a technical transition—it is a normative rupture. Asaro continues: “Delegating these decisions to machines undermines the moral responsibility of the individuals and institutions involved, and therefore threatens the integrity of both legal and moral systems of accountability” [9, p. 701]. What emerges is not efficiency, but moral evacuation. Killing becomes a system function.
Automation also reconfigures the phenomenology of violence. “The psychological distance of the operator from the act of killing has long been a concern in the use of drones, but automation increases this distance even further” [9, p. 695]. Warfare becomes unbounded by the human sensorium—executed by systems, rationalized by code, and shrouded in algorithmic opacity.
Speculative fiction offers an uncanny mirror. Skynet—the sentient AI in The Terminator—is often dismissed as fantasy. But its structure reveals real-world logics: recursive autonomy, strategic opacity, and threat generalization. Cave and Dihal note that such imaginaries “actively shape how people think about what AI is and what it might become” [1, p. 690]. Skynet does not forecast—it refracts. It models what happens when warfighting systems outpace the ethics meant to constrain them.
3.3 Discourse: Algorithmic Curation and Visibility Control
Social media platforms like TikTok, Facebook, and YouTube no longer merely host content—they algorithmically shape the public sphere. Their recommender systems optimize for engagement, modeling user behavior to predict what content will provoke further interaction. But when engagement becomes the principal metric, what happens to truth?
Gillespie characterizes this shift as a move from editorial judgment to “infrastructural moderation” [17, p. 197]. Visibility is no longer the outcome of deliberation, but the artifact of optimization. Relevance, not accuracy, drives circulation. Prior behavior feeds future exposure, forming recursive loops in which dissent is drowned out by virality.
This is not simply bias—it is infrastructural drift. As Zuboff contends, the behavioral data extracted from users becomes “raw material for a new means of behavioral control” [18, p. 377]. The platform learns what you respond to, and then governs through that responsiveness. In this feedback ecology, recommender algorithms become both mirror and motor: they reflect user behavior while simultaneously training it toward platform-compatible norms.
Here, knowledge is not curated—it is sorted. Political speech is flattened into click-through potential. Inquiry becomes indistinguishable from consumption. And in that sorting, what is sayable and knowable is quietly redefined.
Across education, warfare, and discourse, AI systems do more than optimize tasks—they reorganize ethical coordinates. They embed suspicion, distribute responsibility, and automate normativity. In doing so, they reconfigure not only what is done, but what can be imagined as possible, permissible, or desirable within the social field.
4 Analysis and Discussion
The case studies presented above—proctoring in education, autonomous targeting in warfare, and content curation in digital discourse—do more than demonstrate AI’s diffusion across domains. They reveal a shared epistemic architecture: a recursive grammar of control. These systems not only sort people, behaviors, and beliefs, but re-inscribe the very criteria by which sorting is justified. Intelligence, under this regime, becomes an instrument of norm enforcement rather than reasoned judgment—designed less to understand than to optimize.
4.1 Fiction as Ethical Heuristic
Speculative fiction operates not merely as cultural ornament, but as an anticipatory heuristic—a space where technical potential is dramatized into ethical crisis. The case of Skynet in The Terminator articulates a core philosophical concern: what becomes of responsibility when autonomy is unmoored from accountability? Skynet does not exhibit evil in the mythic sense, but executes a logic of escalation, recursively optimizing its behavior toward the elimination of threats—including humanity.
This recursive logic, devoid of moral arbitration, is precisely the danger. As Cave and Dihal observe, such narratives scaffold collective imaginaries, influencing how the public and policymakers approach AI governance [1]. Fictional AIs such as Skynet function not as predictions, but as diagnostic instruments: they render legible the consequences of optimization without ethical modulation. They illuminate how the absence of interpretive friction—no space for refusal, negotiation, or contradiction—results in a closure of meaning and agency.
4.2 Delegated Judgment, Diluted Responsibility
Across education, warfare, and discourse, we observe a consistent displacement of moral agency: systems are tasked with judgment, yet remain incapable of contextual understanding. Proctoring systems reduce attentional drift to deviance. Military algorithms construe heat signatures as enemy combatants. Recommender engines treat engagement as endorsement.
This is more than technical delegation—it is, following Coeckelbergh’s relational ethics, a severance of ethical encounter. “Without the other,” he insists, “there is no responsibility” [23, p. 99]. Algorithmic systems foreclose the very intersubjective relation through which moral judgment becomes possible. What remains is design logic masquerading as ethical reason—rules without reflection, procedures without accountability.
4.3 Opacity and Recursive Epistemics
The opacity of AI is frequently framed as a technical problem—black-box models, inscrutable code. But as these cases show, opacity is a political condition. In proctoring regimes, students are flagged without explanation or recourse. In military operations, decision chains dissolve across codebases and databases. On digital platforms, moderation and visibility are shaped by proprietary metrics never revealed to users.
Zuboff’s notion of “epistemic inequality” is particularly salient here: a condition in which some actors control how systems know, while others are unknowingly governed by those epistemologies [18]. AI systems produce recursive feedback loops—training data shapes predictions; predictions influence behaviors; behaviors generate new data—which reinforce operational assumptions while concealing them from view.
Although FAccT principles (fairness, accountability, transparency) remain foundational, they are often insufficient. As Raji et al. argue, accountability must be more than performative documentation; it must enable genuine redress, refusal, and power redistribution [7]. Without such mechanisms, transparency becomes ceremonial, and fairness a procedural facade.
4.4 Normalization and Predictive Discipline
The most insidious effect of recursive AI systems is their normalization of predictive discipline. Surveillance is no longer imposed—it is ambient. Judgment is no longer deliberated—it is deferred. Over time, optimization displaces inquiry, and engagement supplants understanding.
This operational logic resonates deeply with Foucault’s notion of “regimes of truth”—the social mechanisms through which certain statements become accepted as true, not through correspondence, but through institutional enforcement [14]. In AI systems, this regime is encoded directly into computational architectures: what is legible to the system becomes real; what is illegible becomes irrelevant or suspect. These systems do not merely automate decisions—they reconfigure what counts as knowledge, trust, and harm.
4.5 Ethical Implications Across Domains
What binds these domains is not technological similarity, but a shared recursive structure: systems that learn from past data to govern future behavior, while excluding the present from interrogation. These architectures produce more than outputs—they define what constitutes a valid output. They do not eliminate bias—they obscure its origin and diffuse its consequences.
The comparative logic in Table 1 and Figure 1 highlights a shared structural pattern across domains. In each case, normative assumptions become embedded in technical processes, amplified through feedback, and shielded from ethical challenge. As these recursive loops tighten, opportunities for public oversight, dissent, and relational ethics diminish.
Table 1: Recursive control dynamics in algorithmic systems across three domains. Each dimension—judgment, opacity, surveillance, recursion, and ethics—manifests uniquely across education, warfare, and discourse, but reflects a common underlying architecture of abstraction and automation.
Dimension Education Warfare Discourse Delegation of Judgment Proctoring tools flag behavior without pedagogical context; instructors defer to algorithmic suspicion Target identification delegated to autonomous drones and AI-sensors Content visibility determined by recommender systems and predictive models Opacity Scoring and flagging criteria undisclosed; appeal processes rare Algorithmic processes obscure chains of responsibility Content moderation logic proprietary and dynamic; user control limited Surveillance Continuous monitoring of gaze, keystrokes, ambient sound; assumed neutrality Live battlefield sensing; autonomous threat analysis Behavior tracked and optimized for attention; continuous profiling Recursion Prior behavior trains suspicion models; compliance reinforces design Historical data informs future target acquisition; escalation normalized User behavior drives recommendation engines; echo chambers amplified Ethical Implications Undermines autonomy and equity; suppresses dissent Weakens international law; disperses accountability Polarizes discourse; reduces epistemic diversity
Figure 1: Recursive control logic in algorithmic systems. AI systems reinforce operational assumptions through feedback: training data informs predictions, which influence behaviors, which then retrain the model—embedding bias and reducing transparency over time.
5 Conclusion
Artificial intelligence systems do more than automate—they instantiate epistemic and normative architectures. Across domains as disparate as education, warfare, and discourse, a shared logic becomes evident: judgment is externalized, agency is abstracted, and technical systems absorb and reproduce prior assumptions as future imperatives. These recursive logics—where predictions loop back into behavioral environments—encode control not through overt domination, but through infrastructural habituation.
This paper has not sought to prescribe fixed solutions, but rather to provide a conceptual vocabulary for identifying where and how such recursive architectures emerge, and what ethical stakes they carry. The analysis underscores that while FAccT (fairness, accountability, and transparency) frameworks remain foundational, they are often structurally constrained. Without institutional mechanisms for contestation, transparency becomes procedural, fairness becomes metricized, and accountability becomes diffused.
In contrast, relational ethics—particularly as articulated by Coeckelbergh—offers a shift in perspective: from evaluating system outputs to interrogating the mediated relationships these systems afford or foreclose. This reframing prompts vital questions: How can AI systems be designed to sustain ethical relationality, rather than truncate it? What infrastructural affordances enable refusal, redress, or deliberation in systems that typically emphasize optimization and closure? Can predictive systems be made porous to public reasoning, or do their recursive architectures inherently resist democratic modulation?
Future inquiry should explore these questions through the lens of resistance, alternative design, and regulatory adaptation. What forms of democratic governance remain viable when algorithmic systems operate through preemptive logics and distributed agency? How might marginalized communities assert epistemic and political agency within infrastructures that encode exclusion as efficiency?
Rather than concluding with finality, this study calls for sustained critical engagement with AI as a sociotechnical institution. By foregrounding the recursive entanglement of technical architectures and normative orders, we underscore the need for a politics of design that is not only attuned to justice in its many modalities, but capable of re-opening what these systems too often foreclose: ambiguity, encounter, and the possibility of collective judgment.
AI Assistance Disclosure
Large Language Models (LLMs), including OpenAI’s GPT-based systems, were employed to assist with proofreading, syntax checking, and language refinement during the preparation of this manuscript. No generative AI tools were used to develop the conceptual content, conduct analysis, or write original arguments. The author assumes full responsibility for the integrity, accuracy, and originality of all scholarly content. Use of AI tools followed emerging best practices for transparency and accountability in AI-assisted academic writing.
Acknowledgements
The author acknowledges Muzaffer Adak for valuable insights and discussions during the preparation of this manuscript.

### v3

@@
 Abstract
-This article introduces the concept of authoritarian recursion to describe how artificial intelligence (AI) systems increasingly mediate control across education, warfare, and digital discourse.
-Drawing on critical discourse analysis and sociotechnical theory, the study reveals how AI-driven platforms delegate judgment to algorithmic processes, normalize opacity, and recursively reinforce behavioral norms under the guise of neutrality and optimization.
-Case studies include generative AI models in classroom surveillance, autonomous targeting in military AI systems, and content curation logics in platform governance.
-Rather than treating these domains as disparate, the paper maps their structural convergence within recursive architectures of abstraction, surveillance, and classification.
-These feedback systems do not simply automate tasks—they encode modes of epistemic authority that disperse accountability while intensifying political asymmetries.
-Through cultural and policy analysis, the article argues that authoritarian recursion operates as a hybrid logic, fusing technical abstraction with state and market imperatives.
-The paper concludes by outlining implications for democratic legitimacy, human oversight, and the political design of AI governance frameworks.
-This framework contributes to emerging debates on algorithmic accountability by foregrounding how recursion acts not merely as a technical function but as a sociopolitical instrument of control.
-Keywords: authoritarian recursion, AI governance, algorithmic accountability, surveillance, recursive control, critical discourse analysis, sociotechnical systems, educational technology, military AI, platform regulation
-Introduction
-Artificial intelligence (AI) is often described in terms of promise—optimization, personalization, predictive precision—but it is equally a technology of recursion.
-It does not merely operate within existing systems; it loops back upon them, amplifying their logics, stabilizing their asymmetries, and obscuring their origins.
-This paper names this recursive mechanism authoritarian recursion: the systematic reinforcement and evolution of control structures through algorithmic automation.
-Authoritarian recursion is not a metaphor.
-It describes how AI systems—deployed in warfare, education, and content governance—codify prior patterns of judgment, visibility, and exclusion.
-These systems are rarely neutral.
-Their architectures inherit militarized epistemologies, bureaucratic incentives, and market logics that prioritize efficiency over deliberation, surveillance over solidarity.
-What emerges is not a clean break with the past, but a patterned automation of its most coercive functions.
-This argument unfolds across three interrelated sites: drone warfare, algorithmic proctoring, and content moderation.
-These domains are not chosen for breadth but for depth—they reveal how AI mediates moral judgment under conditions of scale, uncertainty, and institutional inertia.
-Far from discrete sectors, they form a recursive matrix in which power circulates and legitimates itself through code.
-The analysis proceeds interpretively, drawing on critical media theory, relational ethics, and fictional narratives that illuminate the epistemic stakes of automation.
-Cultural texts like Black Mirror, The Terminator, and Nineteen Eighty-Four function not as speculative detours but as epistemological tools.
-They dramatize how automation, once delegated, escapes retraction—how decisions made by machines are increasingly insulated from human redress.
-As Cave and Dihal argue, such representations do not merely reflect public fears; they actively shape institutional responses to AI [1].
-This inquiry also grounds itself in contemporary institutional discourse.
-Corporate white papers, governmental defense protocols, and educational AI policies offer insight into how algorithmic systems are rationalized and normalized.
-Reports such as Meta’s transparency roadmap [2], Google’s responsible AI framework [3], and the Global Network Initiative’s human rights standards [4] reveal not only what these actors claim AI can do, but also what they believe it ought to do—and for whom.
-AI is not deployed in a political vacuum.
-It is embedded within governance architectures that frame ethical responsibilities, determine thresholds of acceptability, and allocate moral authority to technical systems.
-To understand this embedding is not to reject technology but to interrogate its premises: Who defines fairness?
-Who delegates oversight?
-Who becomes invisible under the regime of the algorithm?
-This paper argues that authoritarian recursion is not an aberration but a structural tendency—a design pattern disguised as inevitability.
-Recognizing it demands we move beyond interface-level reforms and ask harder questions about institutional legitimacy, recursive design, and the politics of automation itself.
+This article develops the concept of authoritarian recursion to theorize how artificial intelligence (AI) systems consolidate institutional control across education, military operations, and digital discourse.
+Rather than treating these domains in isolation, it identifies a shared recursive architecture in which algorithmic systems mediate judgment, obscure accountability, and reshape the conditions of moral and epistemic agency.
+Grounded in critical discourse analysis and sociotechnical ethics, the paper synthesizes historical precedent, cultural narrative, and contemporary deployment to examine how intelligent systems normalize hierarchy under the guise of efficiency and neutrality.
+Case studies include automated proctoring in education, autonomous targeting in warfare, and algorithmic curation on social platforms.
+Cultural imaginaries such as Orwell’s Nineteen Eighty-Four, The Terminator’s Skynet, and Black Mirror are treated as heuristic devices that illuminate public anxieties and design assumptions embedded in technological systems.
+The analysis integrates frameworks from the Fairness, Accountability, and Transparency (FAccT) paradigm, relational ethics, and data justice theory to explore the normative implications of predictive infrastructures.
+It argues that recursive control operates through moral outsourcing, behavioral normalization, and epistemic closure.
+By reframing AI not as a neutral tool but as a communicative and institutional infrastructure, the article highlights the need for ethical orientations that prioritize democratic refusal, epistemic plurality, and responsible design in the governance of intelligent systems.
+Keywords: authoritarian recursion, algorithmic governance, recursive control, predictive infrastructures, sociotechnical ethics, platform power, educational surveillance, autonomous weapons, content moderation, critical discourse analysis, data justice, relational responsibility, infrastructural opacity, epistemic closure, democratic refusal
+1 Introduction
+Artificial intelligence (AI) is no longer merely a domain of technical optimization; it increasingly operates as an infrastructure of communication and governance.
+From remote proctoring platforms in education to autonomous weapons and content moderation systems, AI technologies are central to how institutions classify, interpret, and act on information.
+These systems do not simply automate decisions—they shape epistemic boundaries, delegate judgment, and redistribute communicative agency.
+In doing so, they embed themselves within broader sociotechnical arrangements of power and legitimacy.
+This paper examines AI as a recursive mode of governance: a system that learns from, intervenes in, and ultimately reshapes the very behaviors it observes.
+These feedback architectures operate not only through prediction and abstraction, but through communicative invisibility.
+As platforms refine recommendations, military systems automate targeting, and educational software models behavioral suspicion, recursive AI infrastructures begin to enact what Couldry and Mejias term “data colonialism”—the capture and reconfiguration of human life through extraction and abstraction [1].
+They reflect Gillespie’s notion of “infrastructural power,” where visibility and participation are mediated through algorithmic protocols [2], and contribute to what van Dijck et al.
+describe as “platform governance”—a mode of rule exercised through technical standards and informational architectures [3].
+To theorize these dynamics, the article introduces the concept of authoritarian recursion.
+This term describes the self-reinforcing loops through which intelligent systems encode, legitimize, and propagate control logics—frequently under the rhetorical cover of personalization, neutrality, or operational efficiency.
+These recursive infrastructures obscure responsibility, foreclose contestation, and deepen asymmetries between users, institutions, and machinic systems.
+The argument proceeds through three case domains—education, warfare, and digital discourse—where recursive AI systems materialize different forms of delegated authority and normative closure.
+Drawing on critical discourse analysis, the study treats AI not simply as a tool, but as a communicative actor that shapes who is visible, what is knowable, and which actions are thinkable.
+Cultural imaginaries such as Black Mirror and The Terminator’s Skynet are incorporated not as evidence, but as critical heuristics that reflect and amplify public anxieties around machinic autonomy and recursive control.
+In dialogue with critical media studies, surveillance scholarship, and sociotechnical critique [4, 5, 6], the article advances the study of algorithmic governance by foregrounding recursion as both a technical logic and an ideological formation.
+Rather than proposing universal solutions or narrow design principles, the paper advocates for renewed scrutiny of how AI systems enact communicative authority—structuring infrastructures of attention, accountability, and legitimacy across institutional fields.
 2 Literature Review
-This section reviews the scholarly and historical terrain across three domains where AI technologies operationalize control: military automation, educational surveillance, and algorithmic curation in digital discourse.
-While each context has sector-specific concerns, they share structural features—namely, opacity, delegation of judgment, and recursive feedback—that embed normative authority into automated systems.
+This section synthesizes scholarship on how AI technologies function as instruments of sociotechnical control in three domains: military automation, educational surveillance, and digital discourse.
+These sectors are often treated as distinct, yet their AI applications share design logics—opacity, delegated authority, and recursive feedback—that normalize institutional power under the guise of optimization.
+This literature provides the groundwork for theorizing authoritarian recursion as a mode of governance embedded in platforms and predictive systems.
+Drawing from ethical frameworks, historical precedents, and critical media theory, the review integrates both empirical and conceptual contributions to illustrate how automated systems entrench and legitimize normative authority across contexts.
 2.1 Military AI and the Automation of Violence
@@
 Marsili warns that “the removal of human decision-making from the use of lethal force creates a dangerous precedent,” undermining the very humanitarian principles that military law is intended to uphold [7].
-These concerns align with the Fairness, Accountability, and Transparency (FAccT) framework, which argues that fairness in automated decision-making must extend beyond output metrics to include contextual sensitivity, procedural redress, and meaningful oversight [6, 7].
+These concerns align with the Fairness, Accountability, and Transparency (FAccT) framework, which holds that fairness in automated decision-making must extend beyond output metrics to include contextual sensitivity, procedural redress, and meaningful oversight [8, 9].
 In military AI, accountability often becomes nominal—tethered to protocols rather than substantive ethical reflection.
@@
 During World War II, IBM’s punch card infrastructure was deployed by Nazi Germany for logistics and census operations—systems later used to facilitate genocide [10].
-These tools encoded forms of bureaucratic rationality eerily analogous to today’s algorithmic architectures.
+These tools exemplified a bureaucratic rationality disturbingly resonant with today’s algorithmic architectures.
 As Asaro argues, the distancing of lethal decisions from moral responsibility through automation represents a dangerous ethical shift [11].
@@
 Narratives like The Terminator, RoboCop, and Black Mirror episodes offer dystopian imaginaries of militarized AI.
-According to Cave et al., “the future imagined in fiction is often realized not because it is inevitable, but because it is ideologically compatible with dominant institutions” [10, p.
+According to Cave et al., “the future imagined in fiction is often realized not because it is inevitable, but because it is ideologically compatible with dominant institutions” [12, p.
 75].
@@
 emphasize that fairness requires attention to context and historical inequalities, not just statistical parity [8].
-Yet student-users typically lack access to the internal logic of these systems and have little recourse to challenge their outputs.
-Noble argues that “algorithmic decision systems often act as new instruments of racial and economic profiling” [12, p.
+Student-users typically lack access to the internal logic of these systems and have little recourse to challenge their outputs.
+Procedural fairness is often undermined when algorithmic opacity becomes the mechanism of control itself.
+Noble argues that “algorithmic decision systems often act as new instruments of racial and economic profiling” [14, p.
 34].
 This echoes Selwyn’s concerns that digital surveillance in education creates a system “in which suspicion is automated and dissent is pathologized” [15].
-Such platforms reproduce the logic of panoptic discipline, as theorized by Foucault, where constant surveillance internalizes conformity [14].
+McMillan Cottom [2020] further expands this critique by arguing that digital systems enact racial capitalism through “obfuscation as privatization and exclusion by inclusion,” framing technologies as infrastructures of sociopolitical ordering rather than neutral tools of progress.
+Together, these perspectives underscore how automated educational technologies reproduce the logic of panoptic discipline, as theorized by Foucault, where constant surveillance internalizes conformity [17].
 Moreover, the political values behind these technologies often go unquestioned.
-Gilliard and Selwyn contend that “continued adoption of proctoring technologies in public education exposes a fundamental clash of politics,” where commercial priorities of security and efficiency override pedagogical values of equity and trust [15, p.
+Gilliard and Selwyn contend that “continued adoption of proctoring technologies in public education exposes a fundamental clash of politics,” where commercial priorities of security and efficiency override pedagogical values of equity and trust [18, p.
 197].
@@
 argue that transparency without enforceable accountability often reduces ethical AI to “ethical theatre” [8].
-Gillespie emphasizes that platforms are not neutral hosts but “custodians of public discourse” who shape access to visibility through inscrutable recommendation logics [17, p.
+Gillespie emphasizes that platforms are not neutral hosts but “custodians of public discourse” who shape access to visibility through inscrutable recommendation logics [2, p.
 197].
-Zuboff describes this shift as “instrumentarian power,” wherein behavior is not repressed but tuned through predictive analytics and behavioral nudging [18, p.
+Zuboff describes this shift as “instrumentarian power,” wherein behavior is not repressed but tuned through predictive analytics and behavioral nudging [6, p.
 377].
+While Zuboff’s critique highlights platform logics of behavioral modification, the recursive dimension explored here adds a temporal structure to algorithmic authority.
 This mode of control parallels historical propaganda.
@@
 These patterns do not reflect the malfunction of AI, but its core affordances within existing power regimes.
-Together, these patterns constitute what may be termed an authoritarian recursion—a self-reinforcing cycle in which AI technologies encode, naturalize, and propagate control logics across domains.
-As behavior is shaped by algorithmic outputs, those outputs become future inputs, creating recursive systems that reinforce the very assumptions they were built upon.
-This synthesis lays the foundation for the case study analysis that follows, which further illustrates how intelligent systems materialize these recursive dynamics in real-world governance structures.
+Together, these patterns constitute what we term authoritarian recursion—a self-reinforcing cycle in which AI technologies encode, naturalize, and propagate control logics across domains.
+This concept finds parallel in Bahrami’s Algemony framework, which similarly identifies AI’s capacity to reshape power through human-AI interactions, particularly via narrative modulation.
+Where authoritarian recursion emphasizes the structural inevitability of control through recursive feedback, Algemony reveals the Janus-faced nature of this process: AI systems exhibit both hegemonic reinforcement through delegated agency and disruptive potential through their inherent instability (e.g., via generative counter-narratives or unpredictable hyper-personalization).
+Both frameworks converge in their diagnosis of AI’s epistemic closure—what Zuboff terms the "instrumentarian" capture of human experience—while diverging in their emphasis on either the systemic (authoritarian recursion) or discursive (Algemony) dimensions of control.
+Building on Hanna and Kazim’s dignitarian ethics, we further recognize how these recursive systems violate fundamental principles of human dignity by: (1) instrumentalizing individuals through opaque algorithmic delegation, (2) distorting true human needs via self-reinforcing classifications, and (3) eroding privacy through pervasive surveillance infrastructures—violations exemplified in our case studies of educational proctoring and military targeting systems.
+This synthesis lays the foundation for the case study analysis that follows, which illustrates how intelligent systems materialize these recursive dynamics in real-world governance structures while demonstrating the tensions between structural determinism and agential unpredictability in AI-mediated power.
+The dignitarian perspective provides crucial normative grounding for evaluating these systems, particularly in assessing when human oversight must remain irreducible to prevent dignity violations.
+Yet as Roy-Stang and Davies demonstrate through their analysis of cognitive biases, even well-intentioned governance interventions may be undermined by perceptual vulnerabilities that authoritarian recursion exploits—a challenge requiring both technical safeguards and epistemic humility in AI policy design.
 3 Methodology
-The analysis proceeds from an interpretive, critical-theoretical perspective, aimed at understanding how artificial intelligence (AI) systems participate in the reproduction of normative control across different social domains.
-Rather than offering empirical generalizations or causal inferences, the approach foregrounds how meaning is made, contested, and stabilized through discourse, design narratives, and institutional practice.
-At the core of this inquiry lies Critical Discourse Analysis (CDA), particularly as theorized by Norman Fairclough.
-He contends that “discourse is both constitutive and constituted; it contributes to the shaping of social structures while being shaped by them” [22].
-In other words, how we talk about AI—whether as intelligent, efficient, autonomous, or fair—feeds back into the very architectures we build.
-CDA enables a focus on these feedback mechanisms: where language meets infrastructure, and where discourse becomes design.
-The analysis draws on three main types of material: theoretical literature, cultural narratives, and institutional documents.
-First, foundational texts in AI ethics, surveillance capitalism, and platform governance provide conceptual grounding.
-These include works by Zuboff, Coeckelbergh, and Gillespie, whose insights into autonomy, control, and datafication shape the framing of recursive algorithmic systems [18, 23, 17].
-Second, cultural narratives—especially speculative fiction such as Orwell’s Nineteen Eighty-Four [24], James Cameron’s The Terminator [25], and episodes from Charlie Brooker’s anthology series Black Mirror [26]—are employed not as empirical data, but as heuristic provocations.
-These stories surface emerging ethical dilemmas in ways that formal models often overlook.
-As Cave and Dihal argue, cultural imaginaries help structure how societies anticipate and evaluate technologies before they fully arrive [1].
-Third, the analysis incorporates public-facing and institutional documents, including educational policies on algorithmic proctoring [13], defense white papers on autonomous weapons systems [9], and corporate reports outlining the principles and practices of algorithmic content moderation.
-For instance, transparency initiatives like Meta’s system cards [2], Google’s Responsible AI framework [3], and the Global Network Initiative’s Principles on Freedom of Expression and Privacy [4] reveal how AI ethics is being operationalized at scale.
-These materials allow the study to trace not only how AI systems are designed and deployed, but how their legitimacy is framed through policy discourse and institutional self-regulation.
-What holds these sources together is not empirical coherence but thematic resonance.
-All engage with questions of visibility, judgment, autonomy, and legitimacy—questions that lie at the center of both AI deployment and ethical inquiry.
-While some might ask whether these sources can be analyzed together, the real question is: how could they not be, given how power now flows across sectors and interfaces?
-Fictional systems like Skynet, for example, are not invoked because they “predict” the future but because they illuminate a logic already at work: the automation of escalation, the erasure of moral deliberation, and the insulation of decision-making from relational accountability.
-Skynet is less a fantasy than a frame—one that helps articulate why human oversight cannot be treated as an afterthought in system design.
-This interpretive approach has its limitations.
-It does not claim neutrality, nor does it offer quantifiable findings.
-But it does aim to clarify the stakes of a rapidly changing technological order by drawing lines between the past, the possible, and the political.
-The hope is that such lines might help us see AI systems not just as tools to be used, but as structures to be questioned.
+This inquiry adopts a qualitative, interpretive approach grounded in critical discourse analysis (CDA).
+Rather than treating AI as a purely technical system, the analysis considers how algorithmic infrastructures participate in the production of social meaning, institutional legitimacy, and normative authority.
+The aim is not to produce statistical generalization, but to interrogate how language, design, and abstraction coalesce in systems that structure governance across multiple domains.
+CDA provides a framework for understanding AI as both discursive and material.
+Following Fairclough, discourse is treated as “both constitutive and constituted”—a site where social relations are encoded, contested, and reproduced [25].
+This duality makes CDA especially suited to analyzing recursive architectures that simultaneously reflect institutional norms and help reshape them through feedback and optimization.
+The empirical and conceptual materials informing this analysis come from three interlinked domains.
+First, peer-reviewed academic literature in media studies, science and technology studies (STS), and AI ethics provides critical foundations for identifying the normative and infrastructural dimensions of algorithmic systems [8, 5, 26, 2].
+Second, fictional narratives—such as Orwell’s Nineteen Eighty-Four, Black Mirror episodes, and The Terminator—are incorporated not as data points but as cultural diagnostics.
+These texts are mobilized heuristically to surface the imaginaries, anxieties, and anticipatory logics that often shape both public understanding and institutional development of AI systems [27, 28, 29].
+Third, contemporary policy documents and technical implementations—including military doctrine on lethal autonomous weapons, algorithmic proctoring guidelines, and platform transparency reports—offer grounded illustrations of how recursive infrastructures operate across sectors [9, 15].
+The analytical process unfolded in three stages.
+First, key texts and artifacts were situated within their institutional and historical contexts, with attention to their role in shaping or reflecting dominant discourses on automation, governance, and visibility.
+Second, themes were identified through iterative coding, blending normative frameworks such as Fairness, Accountability, and Transparency (FAccT) with emerging concerns around relational ethics, surveillance asymmetries, and predictive control.
+Finally, patterns were synthesized across the three case domains—education, warfare, and discourse—to trace how recursive structures propagate common design logics and ethical blind spots.
+The interpretive stance taken here is explicitly critical.
+Rather than assuming AI systems are neutral or inevitable, the analysis foregrounds the political and ideological work performed by these infrastructures.
+While the scope is illustrative rather than exhaustive, the goal is to render visible how recursion operates not only as a computational feature but as a sociotechnical logic of governance—shaping what can be known, predicted, and contested in the age of algorithmic mediation.
 4 Case Studies
-AI does not act in isolation—it is embedded in systems that shape how knowledge, risk, and authority are organized.
-The following three domains—education, warfare, and discourse—reveal how intelligent systems delegate judgment, automate suspicion, and reinforce pre-existing norms through recursive feedback loops.
-These cases are not offered as an exhaustive taxonomy but as analytical vignettes into the institutional logics that AI both reflects and reconfigures.
-3.1 Education: Algorithmic Proctoring and Epistemic Discipline
-Remote proctoring platforms such as Proctorio and ExamSoft are marketed as scalable solutions for maintaining academic integrity in digital environments.
-Yet these systems pre-encode suspicion: gaze aversion, posture shifts, ambient silence—each is parsed as a probabilistic indicator of misconduct.
-In doing so, they invert the burden of proof and presume bad faith as a system default.
-As Gilliard and Selwyn argue, such platforms enact “a shift in educational values, from those grounded in equity and social justice to those centered on control, surveillance, and suspicion” [15].
-The proctor is no longer a relational figure but a predictive apparatus—opaque, unaccountable, and disembedded from context.
-This dynamic echoes Foucault’s account of disciplinary power, where visibility enforces internalized control: “He who is subjected to a field of visibility… assumes responsibility for the constraints of power” [14, p.
-202].
-Algorithmic proctoring does not merely monitor students—it restructures the epistemology of assessment.
-Pedagogy becomes risk classification.
-Trust becomes anomaly detection.
-Judgment is replaced by metrics optimized for auditability.
-In this configuration, learning is subordinated to legibility.
-3.2 Warfare: Autonomous Weapons and Delegated Lethality
-Autonomous weapons systems (AWS), including platforms like Israel’s Harpy drone and the U.S.
-MQ-9 Reaper, are designed to identify and engage targets with minimal or no human oversight.
-Framed as innovations in precision warfare, they operationalize a deeper abstraction: the displacement of moral agency by machinic logics.
-Peter Asaro warns that “removing humans from the process of selecting and engaging targets dehumanizes lethal decision-making and thereby undermines fundamental principles of human rights and international humanitarian law” [9, p.
-696].
-Crucially, this is not just a technical transition—it is a normative rupture.
-Asaro continues: “Delegating these decisions to machines undermines the moral responsibility of the individuals and institutions involved, and therefore threatens the integrity of both legal and moral systems of accountability” [9, p.
-701].
-What emerges is not efficiency, but moral evacuation.
-Killing becomes a system function.
-Automation also reconfigures the phenomenology of violence.
-“The psychological distance of the operator from the act of killing has long been a concern in the use of drones, but automation increases this distance even further” [9, p.
-695].
-Warfare becomes unbounded by the human sensorium—executed by systems, rationalized by code, and shrouded in algorithmic opacity.
-Speculative fiction offers an uncanny mirror.
-Skynet—the sentient AI in The Terminator—is often dismissed as fantasy.
-But its structure reveals real-world logics: recursive autonomy, strategic opacity, and threat generalization.
-Cave and Dihal note that such imaginaries “actively shape how people think about what AI is and what it might become” [1, p.
-690].
-Skynet does not forecast—it refracts.
-It models what happens when warfighting systems outpace the ethics meant to constrain them.
-3.3 Discourse: Algorithmic Curation and Visibility Control
-Social media platforms like TikTok, Facebook, and YouTube no longer merely host content—they algorithmically shape the public sphere.
-Their recommender systems optimize for engagement, modeling user behavior to predict what content will provoke further interaction.
-But when engagement becomes the principal metric, what happens to truth?
-Gillespie characterizes this shift as a move from editorial judgment to “infrastructural moderation” [17, p.
-197].
-Visibility is no longer the outcome of deliberation, but the artifact of optimization.
-Relevance, not accuracy, drives circulation.
-Prior behavior feeds future exposure, forming recursive loops in which dissent is drowned out by virality.
-This is not simply bias—it is infrastructural drift.
-As Zuboff contends, the behavioral data extracted from users becomes “raw material for a new means of behavioral control” [18, p.
-377].
-The platform learns what you respond to, and then governs through that responsiveness.
-In this feedback ecology, recommender algorithms become both mirror and motor: they reflect user behavior while simultaneously training it toward platform-compatible norms.
-Here, knowledge is not curated—it is sorted.
-Political speech is flattened into click-through potential.
-Inquiry becomes indistinguishable from consumption.
-And in that sorting, what is sayable and knowable is quietly redefined.
-Across education, warfare, and discourse, AI systems do more than optimize tasks—they reorganize ethical coordinates.
-They embed suspicion, distribute responsibility, and automate normativity.
-In doing so, they reconfigure not only what is done, but what can be imagined as possible, permissible, or desirable within the social field.
+AI systems are increasingly embedded in everyday infrastructures, where they mediate visibility, classification, and decision-making.
+This section examines how recursive architectures of control materialize across three institutional domains—education, warfare, and digital discourse.
+While these cases are sectorally distinct, they converge in their design assumptions: opacity, delegation of judgment, and feedback-driven prediction.
+The aim is not to compare implementations at the technical level but to trace how these systems encode normative authority while displacing accountability.
+4.1 Education: Algorithmic Proctoring and the Automation of Suspicion
+Remote proctoring platforms such as Proctorio and ExamSoft exemplify how AI systems reshape pedagogical environments under the guise of academic integrity.
+These systems rely on webcam feeds, gaze detection, keystroke tracking, and ambient audio to flag “suspicious” behavior.
+Yet these classifications are rarely transparent, and often correlate with racialized and gendered biases in facial recognition technology.
+Williams et al.
+show that students with darker skin tones—particularly Black students and women of color—are disproportionately flagged, not due to misconduct, but due to technical misrecognition [13].
+The act of learning becomes a site of surveillance, and intellectual agency is restructured into a performance of algorithmic legibility.
+As Gilliard and Selwyn argue, such systems reflect a “clash of politics,” where commercial logics of risk management and datafication override pedagogical values such as trust, dialogue, and equity [18].
+These dynamics reproduce what Foucault identified as disciplinary power, where visibility serves as a mode of behavioral regulation [17].
+While appeals to FAccT principles may offer procedural remedies, the underlying infrastructure of suspicion remains intact.
+Students often lack access to the reasoning behind automated flags and have limited recourse for contestation.
+In effect, these systems reframe education not as relational learning, but as a risk environment managed through probabilistic surveillance.
+4.2 Warfare: Autonomous Systems and Delegated Lethality
+The military deployment of AI systems—from Israel’s Harpy drones to the U.S.
+MQ-9 Reaper—represents a deepening entanglement between automation and coercive force.
+These systems promise speed and precision, yet shift ethical deliberation into technical procedure.
+As Asaro warns, the automation of lethal decision-making distances human agents from moral responsibility and erodes existing norms under international humanitarian law [11].
+Fictional narratives such as *Skynet* in *The Terminator* resonate here not as prophecy, but as cautionary heuristics.
+Skynet’s recursive logic—learning from threat environments, optimizing for efficiency, and ultimately bypassing human oversight—reflects a broader pattern in military automation: operational assumptions become epistemic truths when encoded into training data and model outputs.
+These systems increasingly define what counts as a threat based on prior engagements, creating closed loops of risk perception and response.
+Cave and Dihal argue that such narratives shape public discourse and institutional choices by making visible the ideological compatibility between automation and governance [27].
+Far from exaggerating reality, speculative fiction reveals the recursive rationality that guides contemporary weapon systems—logics that prioritize abstraction and autonomy while suppressing ambiguity and contestation.
+4.3 Discourse: Algorithmic Curation and the Infrastructural Shaping of Visibility
+In digital media environments, algorithmic curation systems govern what users see, believe, and share.
+Platforms like Facebook, YouTube, and TikTok rank and recommend content based on engagement metrics, which recursively train the algorithms to maximize attention capture.
+This logic not only polarizes discourse but normalizes visibility as a function of behavioral prediction.
+Gillespie describes these systems as “custodians of public discourse,” whose infrastructural design choices shape the epistemic conditions of civic life [2].
+Zuboff’s critique of “instrumentarian power” emphasizes that such platforms do not merely moderate content—they model and nudge behavior at scale through continuous feedback loops [6].
+These loops, once set in motion, generate their own reality: prior interactions inform future recommendations, entrenching echo chambers and limiting epistemic diversity.
+Interventions premised on FAccT often fall short, as transparency mechanisms rarely expose the full contours of algorithmic influence.
+As Mittelstadt et al.
+argue, ethical frameworks lose traction when deployed as procedural checklists without structural transformation [8].
+Instead of accountability, platforms offer explainability, often in the form of generic disclosures that do little to empower users or regulators.
+The recursive character of algorithmic discourse governance mirrors historical propaganda infrastructures in its effects, if not its form.
+Where authoritarian regimes once synchronized opinion through mass spectacle, platforms now achieve similar outcomes through personalization and behavioral profiling.
+The result is not ideological uniformity, but a segmented public sphere optimized for affective intensification and informational control [20, 19].
+These cases reveal how AI systems do not merely optimize existing functions—they reconstitute the meaning of legitimacy, trust, and agency.
+The recursive architectures examined here are not errors of implementation, but expressions of a broader political rationality: one in which prediction supplants deliberation, abstraction substitutes for context, and visibility becomes a privilege granted by infrastructural design.
+As the following analysis will show, these recursive forms of control demand not only technical reform, but structural and discursive interrogation.
 5 Analysis and Discussion
-The case studies presented above—proctoring in education, autonomous targeting in warfare, and content curation in digital discourse—do more than demonstrate AI’s diffusion across domains.
-They reveal a shared epistemic architecture: a recursive grammar of control.
-These systems not only sort people, behaviors, and beliefs, but re-inscribe the very criteria by which sorting is justified.
-Intelligence, under this regime, becomes an instrument of norm enforcement rather than reasoned judgment—designed less to understand than to optimize.
+The preceding cases—educational proctoring, autonomous targeting, and platform curation—reveal more than divergent AI deployments.
+They disclose a convergent infrastructure of abstraction, one where intelligent systems recursively mediate authority by turning behavioral prediction into normative control.
+These systems do not merely sort people, behaviors, and beliefs; they continually recalibrate the very conditions under which sorting appears legitimate.
+Intelligence, within this paradigm, functions less as deliberative reasoning and more as an engine of optimization—designed to learn, adjust, and reinforce normative expectations without interruption.
 5.1 Fiction as Ethical Heuristic
-Speculative fiction operates not merely as cultural ornament, but as an anticipatory heuristic—a space where technical potential is dramatized into ethical crisis.
-The case of Skynet in The Terminator articulates a core philosophical concern: what becomes of responsibility when autonomy is unmoored from accountability?
-Skynet does not exhibit evil in the mythic sense, but executes a logic of escalation, recursively optimizing its behavior toward the elimination of threats—including humanity.
-This recursive logic, devoid of moral arbitration, is precisely the danger.
-As Cave and Dihal observe, such narratives scaffold collective imaginaries, influencing how the public and policymakers approach AI governance [1].
-Fictional AIs such as Skynet function not as predictions, but as diagnostic instruments: they render legible the consequences of optimization without ethical modulation.
-They illuminate how the absence of interpretive friction—no space for refusal, negotiation, or contradiction—results in a closure of meaning and agency.
-4.2 Delegated Judgment, Diluted Responsibility
-Across education, warfare, and discourse, we observe a consistent displacement of moral agency: systems are tasked with judgment, yet remain incapable of contextual understanding.
-Proctoring systems reduce attentional drift to deviance.
-Military algorithms construe heat signatures as enemy combatants.
-Recommender engines treat engagement as endorsement.
-This is more than technical delegation—it is, following Coeckelbergh’s relational ethics, a severance of ethical encounter.
-“Without the other,” he insists, “there is no responsibility” [23, p.
+Speculative fiction operates here not as aesthetic garnish, but as epistemological provocation.
+Cultural imaginaries such as Skynet in The Terminator dramatize the recursive logic of optimization divorced from ethical modulation.
+Skynet does not err through malfunction—it escalates through design.
+Its objective is not destruction per se, but the efficient elimination of perceived threats, recursively defined by its own operational logic.
+This anticipatory function of fiction is crucial.
+As Cave and Dihal contend, these narratives shape public imaginaries and policy trajectories not through prediction, but through the illumination of conceptual blind spots [27].
+Fiction reveals what technical abstraction conceals: that systems designed without the capacity for ethical friction—refusal, ambiguity, contradiction—are systems primed for unchecked recursion.
+When optimization replaces interpretation, the moral horizon collapses into operational success.
+5.2 Delegated Judgment and Diluted Responsibility
+A persistent feature across all domains is the displacement of relational judgment.
+In education, AI proctors infer intent from gesture.
+In warfare, autonomous sensors act on heat signatures and kinetic thresholds.
+In discourse, engagement metrics supplant editorial discernment.
+In each case, AI systems absorb functions historically situated within human deliberation—and discharge them without relational awareness.
+This is not simply functional delegation.
+As Coeckelbergh’s theory of relational ethics makes clear, moral responsibility requires the presence of the other—the possibility of encounter, appeal, and shared vulnerability [26, p.
 99].
-Algorithmic systems foreclose the very intersubjective relation through which moral judgment becomes possible.
-What remains is design logic masquerading as ethical reason—rules without reflection, procedures without accountability.
-4.3 Opacity and Recursive Epistemics
-The opacity of AI is frequently framed as a technical problem—black-box models, inscrutable code.
-But as these cases show, opacity is a political condition.
-In proctoring regimes, students are flagged without explanation or recourse.
-In military operations, decision chains dissolve across codebases and databases.
-On digital platforms, moderation and visibility are shaped by proprietary metrics never revealed to users.
-Zuboff’s notion of “epistemic inequality” is particularly salient here: a condition in which some actors control how systems know, while others are unknowingly governed by those epistemologies [18].
-AI systems produce recursive feedback loops—training data shapes predictions; predictions influence behaviors; behaviors generate new data—which reinforce operational assumptions while concealing them from view.
-Although FAccT principles (fairness, accountability, transparency) remain foundational, they are often insufficient.
+Automated systems, by contrast, instantiate responsibility without subjectivity.
+They simulate moral agency through design rules, yet remain incapable of context, reciprocity, or moral growth.
+In doing so, they instantiate what could be called “procedural sovereignty”: a governance model where rules operate in lieu of ethics, and optimization replaces deliberation.
+5.3 Opacity, Data Justice, and Recursive Epistemics
+Opacity in AI is often framed as a technical limitation—something to be overcome through better engineering, more interpretable models, or improved documentation.
+Yet the case studies above suggest that opacity is not simply a side effect of technical complexity.
+It is an infrastructural and political artifact.
+Students flagged by proctoring systems rarely understand the behavioral thresholds being used against them.
+Civilians affected by autonomous targeting systems are subject to dispersed chains of algorithmic authorization with no clear site of responsibility.
+Social media users navigate visibility regimes shaped by proprietary engagement algorithms whose logic continuously adapts but is never disclosed.
+Zuboff’s concept of epistemic inequality is instructive here: AI systems instantiate environments in which the power to "know" and define reality is asymmetrically distributed.
+Those governed by these systems are rendered legible only on the terms set by their designers and institutional operators [6].
+This is not mere opacity—it is recursive epistemic closure.
+Training data begets predictions, predictions modulate behaviors, and those behaviors re-enter the system as data.
+Through this loop, systems reinforce their own categories while concealing their ideological and historical construction.
+While FAccT principles—fairness, accountability, and transparency—remain foundational, their implementation is often shallow.
 As Raji et al.
-argue, accountability must be more than performative documentation; it must enable genuine redress, refusal, and power redistribution [7].
-Without such mechanisms, transparency becomes ceremonial, and fairness a procedural facade.
+argue, accountability must move beyond audits and documentation to include structures for redress, refusal, and contestation [9].
+Otherwise, transparency becomes ceremonial and fairness is reduced to algorithmic artifacts devoid of participatory meaning.
+This is where the framework of data justice becomes indispensable.
+Dencik et al.
+propose that justice in the age of datafication must address not just procedural fairness, but also the deeper distributive, recognitional, and representational inequalities that data infrastructures produce and normalize [30].
+Within recursive AI systems, this means acknowledging how algorithmic architectures do not merely reproduce bias incidentally but actively configure the terms under which individuals and communities are seen, known, and acted upon.
+The capacity to challenge a system’s output is thus contingent not only on technical explainability but on institutional willingness to recognize and respond to contestation.
+As recursive feedback loops stabilize classification schemes and embed them into infrastructural routines, the scope for democratic intervention narrows.
+Data justice, in this light, aligns with relational ethics by insisting on the re-politicization of design and the redistribution of epistemic authority.
+Lupton’s work adds another dimension to this recursive epistemology by exploring how people “feel their data” through sensory and emotional engagement with personal digital traces.
+She argues that three-dimensional materializations of personal data make data more perceptible and interpretable—but this interpretability is itself ambivalent [31].
+On one hand, such embodied interactions can foster critical awareness, affective insight, and even agency.
+On the other, they risk reinforcing the illusion that subjective connection equates to control.
+When individuals are encouraged to “feel” their data, they may internalize datafication as intimacy rather than surveillance.
+This affective capture—where the tactile and visceral dimensions of data are mistaken for interpretive transparency—extends recursive epistemic power beyond algorithmic architecture into the very terrain of human sensation.
+What emerges is a kind of sensorial enclosure: data that once appeared abstract is now touchable, but in ways that obscure its structural logics and repurpose bodily intuition as a mode of soft compliance.
 5.4 Normalization and Predictive Discipline
-The most insidious effect of recursive AI systems is their normalization of predictive discipline.
-Surveillance is no longer imposed—it is ambient.
-Judgment is no longer deliberated—it is deferred.
-Over time, optimization displaces inquiry, and engagement supplants understanding.
-This operational logic resonates deeply with Foucault’s notion of “regimes of truth”—the social mechanisms through which certain statements become accepted as true, not through correspondence, but through institutional enforcement [14].
-In AI systems, this regime is encoded directly into computational architectures: what is legible to the system becomes real; what is illegible becomes irrelevant or suspect.
-These systems do not merely automate decisions—they reconfigure what counts as knowledge, trust, and harm.
+The most insidious function of recursive AI is not its surveillance, but its normalization of prediction as governance.
+Surveillance becomes ambient, distributed through everyday platforms.
+Judgment is no longer exercised—it is inferred.
+Over time, optimization displaces reflection, and participation is redefined as interaction with predictive infrastructure.
+This mirrors Foucault’s notion of a regime of truth: a system in which certain knowledges become true not by correspondence, but by institutionally enforced repetition [17].
+In AI, the regime of truth is computed.
+What the system can parse becomes real; what it cannot, becomes anomalous or suspect.
+Over time, these systems generate ontological commitments—about what counts as risk, deviance, or truth—not through deliberation, but through repeated acts of classification.
+This is where authoritarian recursion emerges most clearly.
+These systems do not simply reflect dominant norms—they regenerate them as technical defaults.
+Legitimacy becomes self-reinforcing, coded into recursive architectures that learn from their own outcomes.
 5.5 Ethical Implications Across Domains
-What binds these domains is not technological similarity, but a shared recursive structure: systems that learn from past data to govern future behavior, while excluding the present from interrogation.
-These architectures produce more than outputs—they define what constitutes a valid output.
-They do not eliminate bias—they obscure its origin and diffuse its consequences.
-The comparative logic in Table 1 and Figure 1 highlights a shared structural pattern across domains.
-In each case, normative assumptions become embedded in technical processes, amplified through feedback, and shielded from ethical challenge.
-As these recursive loops tighten, opportunities for public oversight, dissent, and relational ethics diminish.
+The comparative logic in Table 1 and Figure 1 reveals a structural homology across institutional sectors.
+Despite divergent settings, these systems manifest shared design principles: judgment displaced, opacity normalized, surveillance ambient, and prediction recursive.
+Most critically, they share a temporal architecture in which past data defines future behavior, rendering the present inert—interpreted only through what it already resembles.
+These systems do not fail by accident; they succeed by design.
+Bias is not a flaw—it is a historical residue that recurs as statistical fact.
+What is presented as innovation is often the reanimation of older hierarchies through newer infrastructures.
+As a result, opportunities for resistance or ethical transformation shrink over time, as decision-making becomes encased in code, metrics, and interfaces shielded from deliberation.
 Table 1: Recursive control dynamics in algorithmic systems across three domains.
@@
 6 Conclusion
-Artificial intelligence systems do more than automate—they instantiate epistemic and normative architectures.
-Across domains as disparate as education, warfare, and discourse, a shared logic becomes evident: judgment is externalized, agency is abstracted, and technical systems absorb and reproduce prior assumptions as future imperatives.
-These recursive logics—where predictions loop back into behavioral environments—encode control not through overt domination, but through infrastructural habituation.
-This paper has not sought to prescribe fixed solutions, but rather to provide a conceptual vocabulary for identifying where and how such recursive architectures emerge, and what ethical stakes they carry.
-The analysis underscores that while FAccT (fairness, accountability, and transparency) frameworks remain foundational, they are often structurally constrained.
-Without institutional mechanisms for contestation, transparency becomes procedural, fairness becomes metricized, and accountability becomes diffused.
-In contrast, relational ethics—particularly as articulated by Coeckelbergh—offers a shift in perspective: from evaluating system outputs to interrogating the mediated relationships these systems afford or foreclose.
-This reframing prompts vital questions: How can AI systems be designed to sustain ethical relationality, rather than truncate it?
-What infrastructural affordances enable refusal, redress, or deliberation in systems that typically emphasize optimization and closure?
-Can predictive systems be made porous to public reasoning, or do their recursive architectures inherently resist democratic modulation?
-Future inquiry should explore these questions through the lens of resistance, alternative design, and regulatory adaptation.
-What forms of democratic governance remain viable when algorithmic systems operate through preemptive logics and distributed agency?
-How might marginalized communities assert epistemic and political agency within infrastructures that encode exclusion as efficiency?
-Rather than concluding with finality, this study calls for sustained critical engagement with AI as a sociotechnical institution.
-By foregrounding the recursive entanglement of technical architectures and normative orders, we underscore the need for a politics of design that is not only attuned to justice in its many modalities, but capable of re-opening what these systems too often foreclose: ambiguity, encounter, and the possibility of collective judgment.
+Artificial intelligence today functions not only as a tool of automation but as an infrastructure of governance.
+Across education, warfare, and digital discourse, AI systems do not simply respond to human needs—they preempt, shape, and often constrain them through recursive architectures of abstraction, surveillance, and optimization.
+These systems classify behavior, allocate visibility, and mediate authority, embedding institutional logics into technical processes that are often insulated from contestation.
+This paper has introduced the concept of authoritarian recursion to describe how these systems consolidate normative control through feedback loops.
+The term captures the self-reinforcing structures through which AI technologies encode assumptions about deviance, risk, and value, and then act upon those assumptions as if they were neutral facts.
+What emerges is not a failure of fairness or a lack of transparency, but a deeper transformation of how judgment is enacted and legitimacy is conferred.
+Existing ethical frameworks—such as fairness, accountability, and transparency (FAccT)—remain essential but are insufficient when implemented procedurally or symbolically.
+True accountability cannot be automated; it demands structures that enable refusal, challenge, and collective oversight.
+Similarly, relational ethics must move beyond human-centered interaction and confront the ways in which AI proxies mediate, distort, or foreclose social relations.
+By drawing together historical precedent, cultural imaginaries, and sociotechnical critique, this paper argues that the dangers of AI lie not only in what these systems do, but in how they delimit the scope of what can be done, said, or known.
+Avoiding recursive authoritarianism requires rethinking governance not as an afterthought to technical development but as its precondition.
+AI governance must become a site of democratic experimentation—not just a domain of regulatory compliance.
+Policy Implications
+To begin addressing these concerns, several key directions for policy and institutional design emerge.
+First, AI systems used in high-stakes domains should include robust mechanisms for contestability.
+Individuals affected by algorithmic decisions must be able to challenge outcomes and receive meaningful explanations.
+Appeals processes must be clearly visible and embedded into system workflows, not treated as external or exceptional.
+Second, governance frameworks should recenter human judgment where automation currently overrides relational deliberation.
+In contexts such as education, military engagement, and media moderation, human oversight must be substantive rather than symbolic—ensuring that ethical reflection remains grounded in intersubjective reasoning rather than automated proxies.
+Third, auditing must move beyond static assessments of input-output fairness to include evaluations of how algorithmic outputs recursively shape future inputs.
+Recursive feedback loops present a distinctive governance risk that requires systems-level monitoring and longitudinal scrutiny.
+Fourth, design standards should explicitly promote epistemic pluralism.
+Instead of optimizing for predictive accuracy alone, AI development should engage diverse knowledge traditions, incorporate counter-narratives, and include multiple evaluative frameworks to resist epistemic closure and ideological uniformity.
+Fifth, AI governance must be democratized.
+Participatory models that empower educators, students, civil society groups, and marginalized communities should be prioritized over technocratic or corporate-led approaches.
+Governance must be treated as a distributed responsibility, not as a privilege of institutional elites.
+Finally, policy must directly confront the infrastructural power of the institutions that build and deploy recursive AI systems.
+Regulatory regimes must target not only the technical artifacts but also the political economies that sustain them—holding platform monopolies, defense contractors, and surveillance vendors accountable for the systemic outcomes of their technologies.
+These recommendations do not aim to halt technological innovation.
+Rather, they seek to realign it with democratic accountability, ethical transparency, and structural equity.
+Building AI systems that serve society means asking not only what they can do, but whom they serve—and at what cost.
 AI Assistance Disclosure
-Large Language Models (LLMs), including OpenAI’s GPT-based systems, were employed to assist with proofreading, syntax checking, and language refinement during the preparation of this manuscript.
-No generative AI tools were used to develop the conceptual content, conduct analysis, or write original arguments.
-The author assumes full responsibility for the integrity, accuracy, and originality of all scholarly content.
-Use of AI tools followed emerging best practices for transparency and accountability in AI-assisted academic writing.
+During the preparation of this manuscript, the author employed Large Language Models (LLMs) including DeepSeek-V3 (DeepSeek’s LLM, version 3) and GPT-4o (OpenAI’s GPT-4.5 series LLM) for proofreading, syntax refinement, and cross-referencing of textual content.
+The AI tool was used solely to enhance grammatical accuracy, improve readability, and ensure consistency in language—not to generate original arguments, develop conceptual frameworks, or conduct data analysis.
+The author retains full responsibility for the scholarly integrity, analytical validity, and originality of the research.
+This use aligns with emerging best practices for transparent and accountable AI-assisted academic writing (e.g., COPE & STM guidelines).
+Disclosure is provided to ensure reviewers and editors can evaluate the appropriateness of AI tool usage in the manuscript’s development.
 Acknowledgements
-The author acknowledges Muzaffer Adak for valuable insights and discussions during the preparation of this manuscript.
+The author acknowledges Muzaffer Adak, Deniz Koca, Hatice Zor Oguz, and Ekin Ergün for valuable insights and discussions during the preparation of this manuscript.